The Pythagorean Theorem
C² = A² + B²
Leg A = 5 Hypotenuse = 11 Leg B = ?
This was an example on a work sheet my teacher gave us for practice.
They filled in the numbers into the formula and simplified the problem.
11² = 5² + B²
121 = 25 + B² * minus 25 from both sides.
96 = B²
The third step is to take the principal square root.
√96 = √ B²
√96 = B
4√6 = B Add back in original problem.
*I got 9.79 but they gave the answers as 4√6. So my question is how did they get the radical 4√6? My teacher does not want us to give answers in decimal form, but I can’t figure out how they got the radical number.

Answer by kara!(17)

√96=√(16*6)=√16*√6=4√6
Sq(16)is equal to 4, because 4 square is 16